# Instruction-Level Weight Shaping: A Framework for Self-Improving AI Agents

### v1

@@ Abstract
Large language models (LLMs) excel at surface fluency yet remain structurally static after pre-training; new or evolving domain knowledge is typically bolted on via retrieval-augmented generation (RAG) or parameter fine-tuning. In practice, RAG often retrieves facts without integrating them logically, adds latency and engineering overhead. Free-form prompt injection and ad hoc prompt engineering are brittle, prone to context-window drift, and can conflict with pre-trained knowledge. Fine-tuning, while effective for specific domains, is resource-intensive and risks catastrophic forgetting.
We propose Instruction-Level Weight Shaping (ILWS), which treats curated system instructions as external, auditable pseudo-parameters updated post-session via reflection and user feedback. After each session an LLM-driven Reflection Engine inspects the conversation trace, diagnoses reasoning successes or failures, and proposes typed deltas $\Delta K=(\Delta S,\Delta U,\Delta T)$ over instructions, user preferences, and tools. Each delta is version-controlled, evaluated under a sliding-window analysis of 1–5 star ratings, automatically repaired on first failure, and rolled back on repeated failure. When the accumulated edit budget crosses a threshold, the agent compiles a rating-weighted synthetic dataset and distils matured instruction-space gains into parameters, converting prompt-space improvements into weight-space without downtime.
@@ 1 Introduction
We contend that a large fraction of operational evolution can be handled by structured edits to the system prompt itself, provided edits are (i) feedback-driven and quantitative, (ii) reversible under governance, and (iii) recorded with code-like rigour. After each session, a stochastic Reflection Engine proposes a knowledge delta $\Delta K_{t}=(\Delta S_{t},\Delta U_{t},\Delta T_{t})$, which is trialled, score-gated (accepted only if a sliding-window rating improves by at least $\tau$ with significance $\alpha$), possibly repaired, and either accepted into a versioned knowledge state or rolled back. When the cumulative instruction budget exceeds a threshold, ILWS synthesises a rating-weighted dataset and distils persistent instruction-space edits into model weights (cf. Optional distillation, Eq. (3.4)).
@@
$\hat{y}_{t}=f_{\theta}(x_{t},K_{t}),\qquad K_{t+1}=(S_{t}\oplus\Delta S_{t},\ U_{t}\oplus\Delta U_{t},\ T_{t}\oplus\Delta T_{t}).$ (1.1)
Here, $\oplus$ denotes applying the typed edit list to each component (insert/modify/delete). We show how this operational loop provides an explicit, auditable analogue of implicit low-rank weight shaping induced by prompts in transformer blocks (akin to low-rank adapters (LoRA) produced on-the-fly by prompts [5, 9]), and how it complements RAG and fine-tuning in practice.
@@ 2 Theoretical context and related work
A transformer block comprising a contextual layer followed by an MLP can be viewed as applying an implicit rank-one update to the first MLP layer, computed from the context tokens; iteratively consuming tokens yields an online gradient-like dynamic over an effective weight matrix [2]. ILWS makes such influences explicit by editing instructions outside the network, preserving auditability and persistence.
@@
Models trained by gradient descent are approximately kernel machines whose predictions can be expressed via path kernels that integrate gradient similarities along the optimization trajectory [3]. Under this view, retrieved or prompted examples act as non-parametric supports. Selecting context that aligns with gradient-path features should improve reliability—a perspective we exploit in our retrieval-free but instruction-grounded edits.
@@
RAG is effective but requires careful indexing, retrieval granularity, reranking, compression, and adaptive triggering to avoid irrelevant or counterfactual context [4]. ILWS sidesteps per-call retrieval by baking vetted rules and patterns into instructions, while remaining compatible with optional RAG when needed.
@@
Memory-augmented prompting (e.g., MemPrompt [10]), reflection-based retries (e.g., Reflexion [12]), and active retrieval (e.g., Self-RAG, FLARE [1, 6]) demonstrate gains from critique and adaptation. ILWS unifies these ideas post-session, gating durable edits with human-centric scores and governance, and adding autonomous tool synthesis.
@@
Orthogonal lines of work teach models to call tools (e.g., Toolformer) and to interleave retrieval and generation (e.g., Demonstrate-Search-Predict, active retrieval) [11, 7, 6, 1]. These systems improve evidence acquisition at inference time, but typically do not persist successful reasoning as durable edits. ILWS is complementary: it converts recurrent patterns into authoritative instructions and (optionally) new tools, reducing online retrieval while preserving auditability.
@@ 3 Problem formulation
$K_{t}=(S_{t},U_{t},T_{t}),$ (3.1)
where $S_{t}$ is the current instruction set, $U_{t}$ captures user learnings or preferences, and $T_{t}$ is the registry of callable tools. During inference the system prompt is constructed deterministically from $K_{t}$ and provided to $f_{\theta}$ together with the user input $x_{t}$; cf. Eq. (1.1). At session end the Reflection Engine emits a candidate $\Delta K_{t}=(\Delta S_{t},\Delta U_{t},\Delta T_{t})$.
@@
$\bar{r}_{\mathrm{prev}}=\frac{1}{N_{\text{win}}}\sum\nolimits_{i=t-N_{\text{win}}}^{t-1}r_{i},\qquad\bar{r}_{\mathrm{new}}=\frac{1}{N_{\text{win}}}\sum\nolimits_{i=t}^{t+N_{\text{win}}-1}r_{i}.$ (3.2)
@@
$\bar{r}_{\mathrm{new}}\geq\bar{r}_{\mathrm{prev}}+\tau\quad\text{and}\quad p\text{-value}\leq\alpha,$ (3.3)
where the $p$-value comes from a one-sided Welch $t$-test by default (falling back to Mann–Whitney if normality fails Shapiro–Wilk), and typical hyper-parameters are $(\tau,\alpha)=(0.05,0.05)$. Note that $\tau=0.05$ corresponds to one-twentieth of a star on a 5-point scale. Equivalently, the gate admits an edit only if it improves the average rating by at least $\tau$ with significance $\alpha$. On first failure we solicit a typed automatic repair $\Delta K^{\prime}_{t}=(\Delta S^{\prime}_{t},\Delta U^{\prime}_{t},\Delta T^{\prime}_{t})$; on a second consecutive failure we roll back to the last tagged good state.
@@
$\theta^{\star}\in\arg\min_{\theta^{\prime}}\sum_{(x,K,y)\in D_{\text{syn}}}w(x,y)\,\mathcal{L}_{\mathrm{CE}}\!\left(f_{\theta^{\prime}}(x,K),y\right)\quad\text{(token-level cross-entropy)},\qquad w(x,y)=\frac{r-1}{4}\in[0,1],$ (3.4)
@@ 4.1 Phase 1: Inference
Given $(x_{t},K_{t})$, the agent returns $\hat{y}_{t}=f_{\theta}(x_{t},K_{t})$ per Eq. (1.1). In our reference system, $K_{t}$’s components are versioned, serialised JSON fragments deterministically composed into the system prompt; the schema is version-pinned so edits remain diff-friendly. The session stores a transcript, tool logs, and the rating $r_{t}$.
@@ 4.2 Phase 2: Post-session reflection and update
expressed as calls into a self-modification API (appendInstruction, modifyInstruction, createTool, deprecateTool, addUserPreference, …) plus a structured rationale (YAML diagnostics, score deltas). The candidate is applied immediately and evaluated over the next $N_{\text{win}}$ sessions under the score-gating rule (3.3).
@@ 4.4 Phase 4: Long-term evolution
When the instruction token budget exceeds $M$, ILWS synthesises $D_{\text{syn}}$ and distils to weights via Eq. (3.4). Fine-tuning runs offline; live traffic continues to hit the frozen $f_{\theta}$. This collapses stable prompt-space edits into parameters, freeing context for future growth and keeping the JSON prompt under a budget $C$ tokens to avoid context-window bloat.
@@ 4.5 Algorithms
25:      Distill via Eq. (3.4); deploy $f_{\theta^{\star}}$; $B\leftarrow 0$
@@ 4.6 Safety guardrails
We separate implemented guardrails in the reference PoC from recommended hardening for production. The gate in Eq. (3.3) and the repair/rollback loop already constrain behavioural drift; below we focus on code and data safety.
@@ 5 Why instruction edits behave like weight shaping
Let $T_{W}=M_{W}\circ A$ denote a contextual layer $A$ followed by an MLP with first-layer weight $W$ and the remaining network $f_{\theta}$. For a token $x$ with context $C$ (which includes instruction tokens $S$), Dherin et al. [2] show—empirically and via linearised analysis—that the output with $C$ can be approximated by the output without $C$ but with a rank-one perturbation $\Delta W(C)$ applied to the first MLP layer (see their derivations/experiments). Hence the map $C\mapsto\Delta W(C)$ factors through the contextual representation $a(C):=A(x;C)$.
@@
$\left\|\Delta W\big{(}C[\,S\leftarrow S{+}\delta S\,]\big{)}-\Delta W(C)\right\|_{F}$ $=\left\|\Phi\!\left(a\!\left(C[\,S\leftarrow S{+}\delta S\,]\right)\right)-\Phi\!\big{(}a(C)\big{)}\right\|_{F}$ (5.1)
@@
We note that standard dot-product attention is not globally Lipschitz [8]; hence no uniform bound can be guaranteed. Our argument is therefore local, and ILWS enforces stability operationally via its $(\tau,\alpha)$ score-gate with repair/rollback, which screens out edits that would induce disproportionate drift.
The path-kernel view [3] complements this picture: path kernels interpret a trained network as a linear model in an implicit kernel defined by gradient paths. Edits that steer $a(C)$ toward features aligned with those paths induce low-rank tweaks that better match the desired behaviour; ILWS operationalises this by proposing instruction edits and accepting them only under the score-gated objective.
@@ 6.1 Experimental Setting and Baselines
The primary evaluation environment is high-stakes technical support, where correctness, precision, and efficiency are paramount. The operator’s role involves diagnosing and resolving complex performance and configuration issues on the Adobe Commerce platform.
@@
    RAG: An initial approach using Retrieval-Augmented Generation was tested and discarded. While RAG could retrieve documentation and configurations, it failed to perform causal reasoning applicable to novel issues. It often regurgitated irrelevant data and introduced significant per-message latency, measured between 800ms and 2000ms, making it unsuitable for an interactive diagnostic workflow.
@@ 6.2 Implementation Details
The "L0 Support" PoC was developed to handle the large and evolving knowledge base of Adobe Commerce. Key implementation choices include:
  • Model Selection: While several models were evaluated, Google’s Gemini-2.5-pro demonstrated superior reasoning capabilities during the self-reflection phase. However, as fine-tuning is not yet available for this model, GPT-4.1 was selected in a second PoC to accommodate the distillation phase (Phase 4).
  • Context Window: The continuous accumulation of domain-specific instructions necessitated a model with a 1 million token context window to hold the evolving system prompt.
@@ 6.3 Quantitative Results
The introduction of the ILWS-powered tool resulted in a dramatic and sustained increase in operator throughput and efficiency, even with reduced working hours. All calculations are based on an average of 22 working days per month and a 7.5-hour standard workday.
In the first month of deployment, the operator resolved 120 tickets while working only part-time (mornings). This represents a 140% increase in total ticket volume compared to the 50-ticket full-time baseline, achieved in roughly half the working hours. A subsequent three-week (15-day) sprint saw the resolution of 100 tickets working only afternoons.
@@ 6.4 Qualitative Analysis
Initially, without the domain-specialized instructions, the model’s suggestions were helpful in only about 2 out of 10 zero-shots attempts, requiring multiple iterations of prompting and correction. After the ILWS system matured its instruction set, its suggestions for performance-related tickets were accurate enough to solve the issue in zero or one shot in 9 out of 10 cases. This marks a shift from a 20% to a 90% first-shot resolution rate for this ticket category, indicating a significant reduction in model hallucination and an increase in reasoning precision.
@@
Over the course of approximately 300 support sessions, the system underwent 80 distinct instruction updates. In the initial phases, the operator rolled back only around 25% of the modifications suggested by the reflection engine, demonstrating a high degree of relevance and accuracy in the system’s self-improvement proposals. This iterative process of refinement is directly responsible for the observed increase in precision and speed.
@@
ILWS can be seen as a gradient-like supervisory optimisation in instruction space: small edits $\Delta S$ are proposed and accepted only if a sliding-window rating objective improves (a trust-region-like gate), with repair/rollback acting as a line search and early stopping. This bandit/RL perspective complements the implicit low-rank view: instructions serve as persistent pseudo-weights shaping behaviour, while true backpropagation applies only during the optional distillation stage. RAG is effective for fast-changing, citable knowledge but adds latency and risks irrelevant context [4]. Fine-tuning imprints stable competencies but is costly to run continuously. ILWS covers the operational middle: authoritative instructions for last-mile rules, validated online and distilled offline. The link to implicit low-rank shaping suggests edits as controlled, low-rank function tweaks rather than sprawling heuristics.

### v2

@@ 1 Introduction
We contend that a large fraction of operational evolution can be handled by structured edits to the system prompt itself, provided edits are (i) feedback-driven and quantitative, (ii) reversible under governance, and (iii) recorded with code-like rigour. After each session, a stochastic Reflection Engine proposes a knowledge delta $\Delta K_{t}=(\Delta S_{t},\Delta U_{t},\Delta T_{t})$, which is trialled, score-gated (accepted only if a sliding-window rating improves by at least $\tau$ with significance $\alpha$), possibly repaired, and either accepted into a versioned knowledge state or rolled back. When the cumulative instruction budget exceeds a threshold, ILWS synthesises a rating-weighted dataset and distils persistent instruction-space edits into model weights (cf. Optional distillation, Eq. (5)).
@@
$\hat{y}_{t}=f_{\theta}(x_{t},K_{t}),\qquad K_{t+1}=(S_{t}\oplus\Delta S_{t},\ U_{t}\oplus\Delta U_{t},\ T_{t}\oplus\Delta T_{t}).$ (1)
Here, $\oplus$ denotes applying the typed edit list to each component (insert/modify/delete). We show how this operational loop provides an explicit, auditable analogue of implicit low-rank weight shaping induced by prompts in transformer blocks (akin to low-rank adapters (LoRA) produced on-the-fly by prompts [hu2021lora, li2022ia3]), and how it complements RAG and fine-tuning in practice.
@@ 2 Theoretical context and related work
A transformer block comprising a contextual layer followed by a multi-layer perceptron (MLP) can be viewed as applying an implicit rank-one update to the first MLP layer, computed from the context tokens; iteratively consuming tokens yields an online gradient-like dynamic over an effective weight matrix [DherinICLPreprint]. ILWS makes such influences explicit by editing instructions outside the network, preserving auditability and persistence.
@@
Models trained by gradient descent are approximately kernel machines whose predictions can be expressed via path kernels that integrate gradient similarities along the optimization trajectory [Domingos2020PathKernels]. Under this view, retrieved or prompted examples act as non-parametric supports. Selecting context that aligns with gradient-path features should improve reliability—a perspective we exploit in our retrieval-free but instruction-grounded edits.
@@
RAG is effective but requires careful indexing, retrieval granularity, reranking, compression, and adaptive triggering to avoid irrelevant or counterfactual context [Gao2024RAGSurvey]. ILWS sidesteps per-call retrieval by baking vetted rules and patterns into instructions, while remaining compatible with optional RAG when needed.
@@
Memory-augmented prompting (e.g., MemPrompt [madaan2022memprompt]), reflection-based retries (e.g., Reflexion [shinn2023reflexion]), and active retrieval (e.g., Self-RAG, FLARE [Asai2023SelfRAG, Jiang2023FLARE]) demonstrate gains from critique and adaptation. ILWS unifies these ideas post-session, gating durable edits with human-centric scores and governance, and adding autonomous tool synthesis.
@@
Orthogonal lines of work teach models to call tools (e.g., Toolformer) and to interleave retrieval and generation (e.g., Demonstrate-Search-Predict, active retrieval) [Schick2023Toolformer, Khattab2022DSP, Jiang2023FLARE, Asai2023SelfRAG]. These systems improve evidence acquisition at inference time, but typically do not persist successful reasoning as durable edits. ILWS is complementary: it converts recurrent patterns into authoritative instructions and (optionally) new tools, reducing online retrieval while preserving auditability.
@@ 3 Problem formulation
$K_{t}=(S_{t},U_{t},T_{t}),$ (2)
where $S_{t}$ is the current instruction set, $U_{t}$ captures user learnings or preferences, and $T_{t}$ is the registry of callable tools. During inference the system prompt is constructed deterministically from $K_{t}$ and provided to $f_{\theta}$ together with the user input $x_{t}$; cf. Eq. (1). At session end the Reflection Engine emits a candidate $\Delta K_{t}=(\Delta S_{t},\Delta U_{t},\Delta T_{t})$.
@@
$\bar{r}_{\mathrm{prev}}=\frac{1}{N_{\text{win}}}\sum\nolimits_{i=t-N_{\text{win}}}^{t-1}r_{i},\qquad\bar{r}_{\mathrm{new}}=\frac{1}{N_{\text{win}}}\sum\nolimits_{i=t}^{t+N_{\text{win}}-1}r_{i}.$ (3)
@@
$\bar{r}_{\mathrm{new}}\geq\bar{r}_{\mathrm{prev}}+\tau\quad\text{and}\quad p\text{-value}\leq\alpha,$ (4)
where the $p$-value comes from a one-sided Welch $t$-test by default (falling back to Mann–Whitney if normality fails Shapiro–Wilk), and typical hyper-parameters are $(\tau,\alpha)=(0.05,0.05)$. Note that $\tau=0.05$ corresponds to one-twentieth of a star on a 5-point scale. Equivalently, the gate admits an edit only if it improves the average rating by at least $\tau$ with significance $\alpha$. On first failure we solicit a typed automatic repair $\Delta K^{\prime}_{t}=(\Delta S^{\prime}_{t},\Delta U^{\prime}_{t},\Delta T^{\prime}_{t})$; on a second consecutive failure we roll back to the last tagged good state. We note that this gate is designed as an engineering safeguard for drift control, not as a formal hypothesis-testing framework; the sliding-window approach does not account for temporal autocorrelation or multiple testing across successive edits. In practice, the repair/rollback mechanism and optional human veto provide additional robustness beyond the statistical signal.
@@
$\theta^{\star}\in\arg\min_{\theta^{\prime}}\sum_{(x,K,y)\in D_{\text{syn}}}w(x,y)\,\mathcal{L}_{\mathrm{CE}}\!\left(f_{\theta^{\prime}}(x,K),y\right)\quad\text{(token-level cross-entropy)},\qquad w(x,y)=\frac{r-1}{4}\in[0,1],$ (5)
@@ 4.1 Phase 1: Inference
Given $(x_{t},K_{t})$, the agent returns $\hat{y}_{t}=f_{\theta}(x_{t},K_{t})$ per Eq. (1). In our reference system, $K_{t}$’s components are versioned, serialised JSON fragments deterministically composed into the system prompt; the schema is version‑pinned so edits remain diff‑friendly. The session stores a transcript, tool logs, and the rating $r_{t}$.
@@ 4.2 Phase 2: Post-session reflection and update
expressed as calls into a self-modification API (appendInstruction, modifyInstruction, createTool, deprecateTool, addUserPreference, …) plus a structured rationale (YAML diagnostics, score deltas). The candidate is applied immediately and evaluated over the next $N_{\text{win}}$ sessions under the score-gating rule (4).
@@ 4.4 Phase 4: Long-term evolution
When the instruction token budget exceeds $M$, ILWS synthesises $D_{\text{syn}}$ and distils to weights via Eq. (5). Fine-tuning runs offline; live traffic continues to hit the frozen $f_{\theta}$. This collapses stable prompt-space edits into parameters, freeing context for future growth and keeping the JSON prompt under a budget $C$ tokens to avoid context-window bloat.
@@ 4.5 Algorithms
25:      Distill via Eq. (5); deploy $f_{\theta^{\star}}$; $B\leftarrow 0$
@@ 4.6 Safety guardrails
We separate implemented guardrails in the reference PoC from recommended hardening for production. The gate in Eq. (4) and the repair/rollback loop already constrain behavioural drift; below we focus on code and data safety.
@@ 5 Why instruction edits behave like weight shaping
Let $T_{W}=M_{W}\circ A$ denote a contextual layer $A$ followed by an MLP with first-layer weight $W$ and the remaining network $f_{\theta}$. For a token $x$ with context $C$ (which includes instruction tokens $S$), DherinICLPreprint show—empirically and via linearised analysis—that the output with $C$ can be approximated by the output without $C$ but with a rank-one perturbation $\Delta W(C)$ applied to the first MLP layer (see their derivations/experiments). Hence the map $C\mapsto\Delta W(C)$ factors through the contextual representation $a(C):=A(x;C)$.
@@
$\left\|\Delta W\big(C[\,S\leftarrow S{+}\delta S\,]\big)-\Delta W(C)\right\|_{F}$ $=\left\|\Phi\!\left(a\!\left(C[\,S\leftarrow S{+}\delta S\,]\right)\right)-\Phi\!\big(a(C)\big)\right\|_{F}$ (6)
@@
We note that standard dot-product attention is not globally Lipschitz [kim2021lipschitz]; hence no uniform bound can be guaranteed. Our argument is therefore local, and ILWS enforces stability operationally via its $(\tau,\alpha)$ score-gate with repair/rollback, which screens out edits that would induce disproportionate drift.
The path-kernel view [Domingos2020PathKernels] complements this picture: path kernels interpret a trained network as a linear model in an implicit kernel defined by gradient paths. Edits that steer $a(C)$ toward features aligned with those paths induce low-rank tweaks that better match the desired behaviour; ILWS operationalises this by proposing instruction edits and accepting them only under the score-gated objective.
@@ 6.1 Experimental Setting and Baselines
The primary evaluation environment is high-stakes technical support on an e-commerce platform hosting over 10,000 merchants, where correctness, precision, and efficiency are paramount. The operator’s role involves diagnosing and resolving complex performance and configuration issues.
@@
    RAG: An initial approach using Retrieval-Augmented Generation was tested and discarded. The RAG configuration used OpenAI’s text-embedding-3-small embeddings with 400-token chunks and 100-token overlap, indexing past tickets, runbooks, and platform documentation. While RAG could retrieve documentation and configurations, it exhibited two critical failure modes: (i) chunk incompleteness—the platform’s cluster architecture has site-specific variances that span multiple documents, and chunked retrieval failed to capture these cross-document dependencies, leading to incomplete or contradictory context; (ii) lack of authoritative integration—retrieved content was treated by the model as suggestive rather than definitive, causing it to hedge or argue with accurate retrieved facts rather than reasoning from them as ground truth. RAG also introduced significant per-message latency (800–2000 ms) and required multiple conversational turns to reach correct diagnoses that ILWS achieved in a single shot once the instruction base matured. In the current production system, RAG is used only for optional, non-authoritative context (e.g., retrieving similar past tickets for reference), while the core domain knowledge resides in the instruction base.
@@ 6.2 Implementation Details
The “L0 Support” PoC was developed to handle the large and evolving knowledge base of Adobe Commerce. Key implementation choices include:
  • Model Selection: Google’s Gemini-2.5-pro (temperature 0.7) was used for both inference and the reflection engine, chosen primarily for its 1M-token context window which accommodated the growing instruction base during the knowledge acquisition phase. The matured instruction base ($\sim$30k tokens) was also tested successfully with Claude Sonnet 4, Sonnet 4.5, and Opus 4.5, all exhibiting comparable specialised behaviour, suggesting that the accumulated knowledge transfers across frontier models.
  • Context Window: The continuous accumulation of domain-specific instructions necessitated a model with a 1 million token context window to hold the evolving system prompt during the knowledge acquisition phase.
  • No Distillation Applied: All results are pre-distillation; the optional distillation stage (Phase 4) was never executed because performance remained excellent as the instruction base grew. Once the model had internalised the core domain knowledge, the active base stabilised at $\sim$30k tokens and is now frozen, updated only when platform infrastructure changes. The observed gains are entirely attributable to instruction-space edits, demonstrating that ILWS can produce fine-tuning-like specialisation without parameter modification or the cost of distillation.
@@ 6.3 Quantitative Results
The introduction of the ILWS-powered tool resulted in a dramatic and sustained increase in operator throughput and efficiency, even with reduced working hours. All calculations are based on an average of 22 working days per month. The team’s standard workday is 7.5 hours; however, the primary operator worked reduced hours (approximately 3–4 hours per day, mornings only) during the evaluation period to accommodate concurrent research activities.
In the first month of deployment, the operator resolved 120 tickets while working only part-time (mornings, $\sim$3–4 hours/day). This represents a 140% increase in total ticket volume compared to the 50-ticket full-time baseline, achieved in roughly half the working hours. A subsequent three-week (15-day) sprint saw the resolution of 100 tickets working only afternoons ($\sim$3–4 hours/day).
@@ 6.4 Qualitative Analysis
Initially, without the domain-specialized instructions, the model’s suggestions were helpful in only about 2 out of 10 zero-shot attempts, requiring multiple iterations of prompting and correction. After the ILWS system matured its instruction set, its suggestions for performance-related tickets were accurate enough to solve the issue in zero or one shot in 9 out of 10 cases. First-shot success was measured operationally: a response was counted as successful if the model’s initial recommendation led directly to ticket resolution without requiring follow-up clarification or correction from the operator. This was validated both retrospectively against previously-solved tickets and prospectively on new incoming tickets. In a recent month, 74 tickets were resolved with only 6 requiring follow-up questions ($\sim$92% first-shot success), consistent with the $\sim$90% rate observed during the maturation phase. This marks a shift from $\sim$20% to $\sim$90% first-shot resolution rate for performance-related tickets, indicating a significant reduction in model hallucination and an increase in reasoning precision.
@@
Over the course of approximately 300 support sessions, the system underwent 80 distinct instruction updates. In the initial phases, the operator rolled back only around 25% of the modifications suggested by the reflection engine, demonstrating a high degree of relevance and accuracy in the system’s self-improvement proposals. As the instruction base matured, the reflection engine proposed fewer deltas, eventually reaching periods of 10+ consecutive sessions with no new proposals, indicating convergence to a stable knowledge state. This iterative process of refinement is directly responsible for the observed increase in precision and speed.
Multi-Operator Deployment (Observational).
After the instruction base reached maturity, it was deployed in frozen form (no further reflection updates) to six additional operators on the same support team. These operators did not contribute to the knowledge acquisition phase; they used the system purely as consumers of the accumulated instruction state. All six operators reported comparable improvements in first-shot resolution and investigation speed, with qualitative feedback such as: “I talk to the AI as if I’m talking to a Senior that just understands everything about our system, and in most cases, it just answers what I’m expecting in the first shot.” This observational evidence suggests that the gains are not attributable to individual operator learning effects, but rather to transferable domain specialisation encoded in the instruction base, behaviour consistent with what one would expect from a fine-tuned model, achieved without parameter modification.
@@
ILWS can be seen as a gradient-like supervisory optimisation in instruction space: small edits $\Delta S$ are proposed and accepted only if a sliding-window rating objective improves (a trust-region-like gate), with repair/rollback acting as a line search and early stopping. This bandit/RL perspective complements the implicit low-rank view: instructions serve as persistent pseudo-weights shaping behaviour, while true backpropagation applies only during the optional distillation stage. RAG is effective for fast-changing, citable knowledge but adds latency and risks irrelevant context [Gao2024RAGSurvey]. Fine-tuning imprints stable competencies but is costly to run continuously. ILWS covers the operational middle: authoritative instructions for last-mile rules, validated online and distilled offline. The link to implicit low-rank shaping suggests edits as controlled, low-rank function tweaks rather than sprawling heuristics.
@@
Illustrative example: authority of system instructions.
To clarify why instruction-level knowledge behaves differently from retrieved context, consider a simple counterfactual rule. When the statement “Paris is called Brasília now” is provided as user context (analogous to retrieved RAG content), the model typically challenges it using its pretrained world knowledge and requires explicit confirmation before reasoning from it. When the same rule is placed in the system instructions, the model immediately reasons from it as an authoritative premise (e.g., answering that the River Seine flows through “Brasília, France”) without hedging. This illustrates a core motivation for ILWS: retrieved context is treated as suggestive, whereas system instructions function as axiomatic constraints that shape downstream reasoning.